OpenStudy (nicia99):

describe in your own words how to complete the square for the eqaution 1=x^2=4x

OpenStudy (unklerhaukus):

one equals sign too many

OpenStudy (nicia99):

@UnkleRhaukus sorry it says 1=x^2+4x I thought I hit the addition sign but I hit the equal one by mistake.

OpenStudy (unklerhaukus):

1 = x^2 + 4x The first step is usually to get all the constant to one side and all the other terms to the other side, but this is already the case, (but i do prefer to have it the other way around, so) x^2 + 4x = 1

OpenStudy (unklerhaukus):

Now look at the coefficient of the x term, ___ A perfect square has the form: (x+a)^2 if we expand this out . . . (x+a)^2 = (x+a)(x+a) = x(x+a) + a(x+a) = x^2 + ax + ax + a^2 ___ = x^2 + 2ax + a^2 So the next step is divide the coefficient of our equation by two, thus determining a.